OpenStudy (anonymous):

x/2-3=-4 help me solve this thank you

OpenStudy (anonymous):

\[\frac{x}{2}-3=-4 => \frac{x}{2}=-4+3 => x=-1*2 => x=2\]

OpenStudy (anonymous):

x/2=-1 ==>x=-2

OpenStudy (anonymous):

x = -2

OpenStudy (anonymous):

x = -2

OpenStudy (anonymous):

Oops, i missed the signal, the correct answer is x=-2